is there any method of derivation of formula of relation of sum and product of roots such as |alpha – bita| = …..
as explained by vinti mam in the chapter of quadratic equation?

Grade:11

1 Answers

Vijay Mukati
8 years ago
You can use the identity here, (a-b)2= (a+b)2- 4ab. Here, a+b = sum of roots and ab = product of roots.Thanks.